Assume that the following data is for a profit-maximizing manufacturer:

Quantity $Total Cost
0 100
1 140
2 160
3 190
4 240
5 300
6 370
7 450
8 550




a. Does this manufacturer operate in the short-run or in the long-run? Explain.
b. If this manufacturer shuts down what would be the profit or loss?
c. If the price is $75 per unit, what are the profit maximizing output and the level of profit or loss?
d. If the price is $55 per unit, what are the profit maximizing output and the level of profit or loss?
e. What is the value of the break-even price?
1
Expert's answer
2018-10-31T13:52:09-0400

Q TC MC

0 100 -

1 140 40

2 160 20

3 190 30

4 240 50

5 300 60

6 370 70

7 450 80

8 550 100

a. This manufacturer operates in the short-run.

b. This manufacturer shuts down if P < AVC and the loss is higher than his fixed costs.

c. If the price is $75 per unit, then P = MR = MC = 75 and the profit maximizing output is 6 units and the level of profit is TP = TR - TC = 75*6 - 370 = $80.

d. If the price is $55 per unit, then P = MR = MC = 55, the profit maximizing output is 4 units and the level of loss is TP = TR - TC = 55*4 - 240 = -$20.

e. The value of the break-even price is:

P = FC/Q + VC = TC/Q = ATC.

At Q = 4: P = 240/4 = $60.

At Q = 6: P = 370/6 = $61.67.